OpenStudy (anonymous):

15 cards with the number 8 to 22 are put into a box.a card is drawn at random from the boxfind the probability of drawing a card with a number which is not a prime number.

OpenStudy (anonymous):

from 8-22 the prime numbers are 11,13,17,19 hence the probability is 4/15

OpenStudy (anonymous):

prime numbers from 8-22: 11,13,17,19 so there are 4 so u want non prime numbers so 15-4=11 so probability is 11/15

OpenStudy (anonymous):

the probability required is 1-(4/15)=11/15

OpenStudy (anonymous):

11/15